OpenStudy (anonymous):

what ratio is equivalent to 2/3?

OpenStudy (anonymous):

1:1.5 4:6 6:9 12:18 ... etc., etc., etc...

OpenStudy (anonymous):

All elements of the equivalence class 2/3.

OpenStudy (anonymous):

@Herp_Derp i just needa answer lol i know no pre-algebra

OpenStudy (anonymous):

There is an infinite number of equivalent ratios. Here are a few of them: 3 to 4.5 4 to 6 6 to 9 8 to 12 (66 and 2/3) to 100 792 to 1,188 0.02 to 0.03

OpenStudy (anonymous):

@JvBuenconcejo thanks but those aren't the ratios my stuff is its fraction type

OpenStudy (precal):

ratio is another label for fraction take (2/3) and multiply it by one except one will take the form of a number over the same number example (2/3) times (3/3) (2x3)/(3x3)=6/9

OpenStudy (precal):

another example (2/3) times (5/5) (2x5)/(3x5)=6/10

OpenStudy (anonymous):

@precal i dont get it

OpenStudy (precal):

I just describe the procedure on how to create other fractions that represent 2/3